OpenStudy (anonymous):

angles A and B are located in the first quadrant. if sin A= 5/13 and cos B = 3/5,determine the exact value of cos (A+B)

OpenStudy (campbell_st):

well draw the 2 triangles...  find x and y using pythagoras' theorem then you need to know \[\cos(A + B) = \cos(A)\cos(B) - \sin(A)\sin(B)\] so make the ratio substitutions and then calculate an answer

OpenStudy (jdoe0001):

use the pythagorean theorem to find "x" and "y" for each once you find the missing side for each you'd find the cosine and sine for each corresponding

OpenStudy (jdoe0001):

for example, notice campbell_st picture thus we could say \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse}=\cfrac{b}{c} \qquad \qquad % cosine cos(\theta)=\cfrac{adjacent}{hypotenuse}=\cfrac{a}{c} \\ \quad \\ sin(A)=\cfrac{5}{13}\to \cfrac{b}{c}\to \cfrac{b=5}{c=13}\impliedby \textit{find cos(A) first} \\ \quad \\ c^2=a^2+b^2\implies c^2-b^2=a^2\implies \pm\sqrt{c^2-b^2}={\color{blue}{ a}} \\ \quad \\ thus\implies cos(A)=\cfrac{{\color{blue}{ a}}}{c}\) the one thing about the pythagorean theorem is that the root could give us either, positive or negative values so, which one do we use? well, we know that angle A is in the 1st quadrant and in the 1st quadrant, side "x" or the adjacent side, is positive so we use the positive value from the pythagorean theorem

OpenStudy (jdoe0001):

and then you do the same for the missing side on angle B to get the sin(B) for it once you get those two, cos(A) and sin(B) use as campbell_st suggested -> \(\bf cos({\color{brown}{ \alpha}} + {\color{blue}{ \beta}})= cos({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}})- sin({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}})\)